OpenStudy (anonymous):

find the indicated outputs f(x=5xsquare 2- 2x f(0)= f(-1)= f(2)=

OpenStudy (anonymous):

\[f(x)=5x^2-2x\]?

OpenStudy (anonymous):

\[f(0)=5\times 0^2-2\times 0 = 0\]

OpenStudy (anonymous):

\[f(-1)=5\times (-1)^2-2\times -1=5+2=7\]

OpenStudy (anonymous):

same idea for the last one

OpenStudy (anonymous):

?????